OpenStudy (anonymous):

what are the two x roots of the polynomial below? x^2+5x+11

OpenStudy (anonymous):

use the formula :)

OpenStudy (anonymous):

quadratic formula

OpenStudy (campbell_st):

there are both complex... so use the quadratic formula

OpenStudy (anonymous):

im embarrassed to say that I dont remember how to

OpenStudy (campbell_st):

for a quadratic \[ax^2 + bx + c = 0\] use \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = 1, b = 5 and c = 11

OpenStudy (anonymous):

do you remember now? and dont be embarrassed that's y we are here!!!

OpenStudy (anonymous):

kim?

OpenStudy (anonymous):

my problem is no matter how hard I try, I cant process math equations anymore... :(

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

@campbell_st can you help me?

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

ok np :Dso lets look at x^2+5x+11

OpenStudy (anonymous):

or rather 1x^2+5x+11

OpenStudy (anonymous):

Compare: ax^2 + bx + c and 1x^2 + 5x + 11 what do you think the value of a, b and c are? any ideas?